OpenStudy (anonymous):

Solve for x. 4(x-b)=x A. b=4/3x B. b=3/4x C. x=4/3b D. x=3/4b

OpenStudy (anonymous):

medal will be given! :D

OpenStudy (anonymous):

4x - 4b = x -4b = x - 4x -4b = -3x 4b = 3x b = (3x)/4 Your answer is B.

OpenStudy (anonymous):

Solve for y, then find the value of y when given x=-2 6x=7-4y A. -12 B. 7/4 C. 19/4 D. 24

OpenStudy (anonymous):

medal will be givin again

OpenStudy (anonymous):

4y = 7 - 6x y = 7/4 - 3x/2 f(x) = 7/4 - 3x/2 f(-2) = 7/4 - ( (3(-2) /2) f(-2) = 7/4 + 3 f(-2) = 7/4 + 12/4 f(-2) = 19/4 <------- (Your answer is C.)

OpenStudy (anonymous):

Solve for h. A=1/2bh A. 2A/b=h B. 2A=b/h C. h=1/2Ab D. h=2Ab

OpenStudy (anonymous):

medal will be given again

OpenStudy (anonymous):

A = 1/2 b h 2A = b h h = (2A)/b <----- (Your answer is A.)

OpenStudy (anonymous):

Rearranged it's 2A/b = h which is the same from the question wants you to answer it in =)

OpenStudy (anonymous):

Find the width of a rectangle with a perimiter of 90 and a length of 15. A. 90 B. 15 C. 45 D. 30

OpenStudy (anonymous):

thanks scott :)

OpenStudy (anonymous):

that perimeter prob. was the last one.

OpenStudy (anonymous):

medal will be given

OpenStudy (anonymous):

 Lets see. Area = l x b Perameter = 2l+2b Now you now that l = 15... You also know that b = area / 15 Area = 15 x b So b = 15b/15 So b = b.... o.o Duh??? Lol idk; maybe you can use what I've said and get somewhere?? Area = 15b 30 + 2b = 90 b = Area / 15 30 + 2(Area)/15 = 90.... 2(Area)/15 = 60 2 times the area = 60 x 15 Area = 900/2 Area = 450 Stay with me here xD Back to the start Area = l x b 450 = 15 x b b = 450/15 b = 30 You might want to double check this! xD But I'm confident it's okey saying as it works out as a nice round number :3

OpenStudy (anonymous):

Okay thanks!! :D